OpenStudy (anonymous):

A train travels a distance of 1000 km in 31/2 hours 30 minutes .What is the average speed of the train ?

OpenStudy (anonymous):

The right question is that the train travels 1000km 400m in 31/2 hours 30 minutes . What is the average speed of the train ?

OpenStudy (anonymous):

The answer is 2501/40 km/h. Please help me with the solution.

OpenStudy (anonymous):

The answer is 2501/40 km/h. Please help me with the solution.

OpenStudy (anonymous):

The right question is that the train travels 1000km 400m in 31/2 hours 30 minutes . What is the average speed of the train ?

rishavraj (rishavraj):

1000km 400m = 1000.4 km 30 mins = 1/2 hr

rishavraj (rishavraj):

so 31/2 hrs 30 min = \[\frac{ 31 }{ 2 } + \frac{ 1 }{ 2 } = \frac{ 32 }{ 2 } = 16 ~hrs\]

rishavraj (rishavraj):

s = d/2 \[s = \frac{ 1000.4 }{ 16 }\]

OpenStudy (anonymous):

No that's not the correct answer

rishavraj (rishavraj):

\[\frac{ 1000.4 }{ 16 } = \frac{ 2501 }{ 40 }\]

OpenStudy (anonymous):

Tyy

rishavraj (rishavraj):

yw:)))